# A voice for minorities: diversity in approval-based committee elections under incomplete or inaccurate information

### v1

@@ Appendix B Lower bound for the query complexity of non-adaptive algorithms in the incomplete information setting
[…]
@@
(alg selects at least $\mathbb{P}(\operatorname{\textsc{alg}}\text{ selects at least }q\text{ designated candidates})=\sum_{j=q}^{p}\mathbb{P}(\operatorname{\textsc{alg}}\text{ selects }j\text{ designated candidates})$ (24)
$\leq\sum_{j=q}^{p}\sum_{h=0}^{j}\mathbb{P}\left(\{\operatorname{\textsc{alg}}\text{ selects }j-h\text{ designated candidates by chance}\}\bigcap\,\{\operatorname{\textsc{alg}}\text{ covers }h\text{ of the sets }S_{i}\}\right),$
where the inequality is due to the fact that covering a set $S_{i}$ is necessary to distinguish its designated candidate, but might not be sufficient. For $h=0$, we have bounded the probability above as $p/\ell$. Each of the other at most $(p-q)p$ probabilities is bounded from above by the probability of covering one set $S_{i}$, which is at most $\ell!2^{\ell+t}Up^{2}$. If we fix $U=\frac{\delta}{\ell!2^{t+\ell}p^{3}(p-q)}$ this becomes at most $\delta$. The total probability is then at most $\frac{p}{\ell}+\delta$. Recall that we want this probability to be at most $\epsilon$. We can achieve this, for instance, by fixing both terms to be $\epsilon/2$. Note that we choose $\delta$ (as it determines the query complexity), so it suffices to show that $\frac{p}{\ell}\leq\frac{\epsilon}{2}$. Recall that $p=2k-1$, so that this condition in met when $l\geq\frac{k-1}{\epsilon}$.
Recall that the query complexity was $Um^{2}$ and that $U=\frac{\epsilon/2}{\ell!2^{t+\ell}p^{3}(p-q)}$, where $p=(k-1)/2$. We want to denote the scaling of the complexity in terms of $\gamma$ and $\epsilon$. Recall that $p=(k-1)/2$ and $q=\left\lceil\left(1-\frac{1}{\operatorname{\text{e}}}\right)\gamma+\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)2\gamma-1\right)p-1\right\rceil$. We then get
@@
where the last step required $22k\gamma/2\geq 1/$ 2, which is true in case $k\geq 2,\gamma>1/2$. It is not hard to see that $\frac{1}{\ell!2^{t+\ell}p^{3}}\leq 1$, so that the final query complexity becomes $\frac{m^{2}\epsilon}{1-\gamma}$. ∎
Theorem B.1 motivates the study of adaptive algorithms. Ideally, we would have more precise non-asymptotic results, especially because the upper bound of Algorithm 4 contains large constants. Or otherwise, it might be possible to extend this result for a larger exponent of $m$ (as done in [15]), by adapting the instance. Moreover, obviously, a lower bound for adaptive setting is desirable. Both are left for future work.
Appendix C Additions to Section 5
@@
$U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$;
For all $i\in V,\,j\in C$, perform query $(i,j)$ $U$ times and take the majority winner as the ‘true’ query response;
Let $W=[\ ]$ be an empty list;
@@
    For all $c^{\prime}\notin W$, determine $\Delta(W,c^{\prime})$.
@@
    Append $W$ with $\operatorname*{argmax}_{c^{\prime}\notin W}\,\Delta(W,c^{\prime})$.
@@
Theorem 5.1.
In the query model with inaccurate query responses and parameter $p\in(0,\frac{1}{2})$, for any $\delta>0$, $n,m\in\mathbb{N}$, we attain a $(1-1/\operatorname{\text{e}})$-approximate solution to Max Cover with probability at least $1-\delta$ using $nm\cdot U$ queries, where $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$. This gives a query complexity in $\mathcal{O}(nm\log(nm/\delta))$.
@@
We adapt Algorithm 1: for all pairs $(i,j)$ for $i\in V$ and $j\in C$, we perform query $(i,j)$ $U$ times and take the majority winner of the results to be the ‘true’ value, breaking ties arbitrarily. There are $nm$ pairs $(i,j)$ for $i\in V$ and $j\in C$, so we only need to show that for any $\delta>0$, taking $U$ as stated, we obtain an $(1-1/\operatorname{\text{e}})$-approximate solution with probability $1-\delta$.
Take $\Omega=\{0,1\}^{U}$, $\mathcal{F}=\mathcal{P}(\Omega)$ and $\mathbb{P}:\mathcal{F}\to[0,1]$ that satisfies ($\mathbb{P}(\{\mathbf{v}\})=(1-p)^{y}p^{U-y}$ when $\mathbf{v}$ contains $y$ 1s and $U-y$ 0s. Accordingly, write $\mathbf{v}_{i,j}$ for a vector that is the outcome of $U$ throws of $(i,j)$, where entry 0 means the outcome is correct (success) and 1 means the outcome is corrupted (failure). Then take $X:\Omega\to\operatorname{\mathbb{N}}:\mathbf{v}_{i,j}\mapsto\sum_{l=1}^{U}v_{i,j,l}$ for the random variable that maps an outcome vector to the number of successes and write $\bar{X}:=U-X$ for the number of failures.
A Chernoff bound gives $\mathbb{P}(\bar{X}\geq\frac{1}{2}U)\leq\exp(-UD(\frac{1}{2}||p))$ [14] and we can reformulate to
$\begin{split}\exp\left\{-UD\left(\frac{1}{2}||p\right)\right\}&=\exp\left\{-U\left(\frac{1}{2}\log\left(\frac{1/2}{1-p}\right)+\frac{1}{2}\log\left(\frac{1/2}{p}\right)\right)\right\}\\ &=\exp\left\{-U\frac{1}{2}\log\left(\frac{1/2}{1-p}\right)\right\}\exp\left\{-U\frac{1}{2}\log\left(\frac{1/2}{p}\right)\right\}\\ &=\left(\frac{1}{4(1-p)p}\right)^{-\frac{1}{2}U}.\end{split}$ (26)
This gives an upper bound on the probability that we have an incorrect majority for a single $(i,j)$, where each throw is correct with probability $1-p$. We choose $U$ so that this stays smaller than $\frac{\delta}{nm}$, and for that we need
$-\frac{1}{2}U\geq\log_{\frac{1}{4p(1-p)}}\left(\frac{\delta}{nm}\right)=\frac{\log\frac{\delta}{nm}}{\log\frac{1}{4p(1-p)}}=-\frac{\log\frac{nm}{\delta}}{\log\frac{1}{4p(1-p)}},$ (27)
so we can take $U=\lceil-2\frac{\log((nm)/\delta)}{\log(4p(1-p))}\rceil$. Then the probability of any of the $nm$ estimates being incorrect, is bounded above by $\delta$. ∎
Theorem 5.2.
@@
In the query model with inaccurate query responses, for any $\delta\in(0,1),\,n,m\in\mathbb{N}$, any algorithm that outputs an ($1-1/\operatorname{\text{e}}$)-approximate solution to Max Cover with probability at least $1-\delta$, makes $\Omega(nm\log(1/\delta))$ queries in expectation.
@@
To establish the theorem, we adapt the proof of Theorem 9 in [19], using a result from multi-armed bandit theory [20] to lower bound the query complexity. The Kullback-Leibler divergence between two probability distributions is defined as $D(P||Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}$ for discrete probability functions $P$ and $Q$, as long as, for all $x,\,Q(x)=0\implies P(x)=0$ (otherwise we define it as +$\infty$+ ∞). This reduces to $d(a,b)=a\log\frac{a}{b}+(1-a)\log\frac{1-a}{1-b}$ for two Bernoulli distributions with parameters $a,b\in(0,1)$.
@@
[…]
Lemma C.1.
@@
For any $k\geq 2$, consider an instance $\nu$ with $m$ candidates and $n$ voters. We partition the voters into $m-w$ disjoint “parties" $P_{1},\ldots,P_{m-w}$, where $w\geq 0$ and $w$ is bounded from above by some fraction of $m$. Voters in party $P_{i}$ approve precisely $c_{i}$. Candidates $c_{1},\ldots,c_{m-w-k}$ (together $C_{1}$) all have one approving voter. These $m-w-k$ voters (together $V_{1}$) are alone in their party. Candidates $c_{m-w-k+1},\ldots,c_{m-w}$ (together $C_{2}$) all have $u>\frac{1}{1-\frac{1}{\operatorname{\text{e}}}\cdot k}$ approving voters (together $V_{2}$), so that the only $(1-\frac{1}{\operatorname{\text{e}}})$-approximate solution consists of the $k$ candidates in $C_{2}$:
@@
The remaining $w$ candidates (together $C_{3}$) have no approving voters. Finally we have $V_{4}$, a set of voters that approve of no candidate (homeless voters). For any $m,n$, we can find constants $D,E\in(0,1)$ such that $|C_{1}|\geq D\cdot m$ (which means we need $w\leq m-Dm-k$) and $(1-E)n-ku\geq|C_{1}|$ so that $|V_{4}|\geq E\cdot n$, so let’s pick any such constant $D$ and $E$. Now consider the instance $\nu^{\prime}$ that is obtained by taking, $x:=\lceil k(\frac{u}{1-1/\operatorname{\text{e}}}-1)\rceil$ homeless voters and distributing them (as) evenly (as possible) over $k$ candidates $c_{1},\ldots,c_{k}\in C_{1}$. Other than that, all preferences remain unchanged. With this value of $x$, we have $x+k>\frac{ku}{1-1/\operatorname{\text{e}}}$, which means that $\nu^{\prime}$ has a unique $(1-1/\operatorname{\text{e}})$-approximate solution in $c_{1},\ldots,c_{k}\subseteq C_{1}$, whereas the unique $(1-1/\operatorname{\text{e}})$-approximate solution of $\nu$ was the set $C_{2}$. Since both instances have only one optimal solution, and these differ, we have $L_{\nu}\cap L_{\nu^{\prime}}=\emptyset$.
@@
[…]
So now we can rewrite the inequality in Lemma C.1 as
@@
$\sum_{C_{1,k}}\sum_{V_{4,x}}\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\geq\left\lfloor\frac{En}{x}\right\rfloor\left\lfloor\frac{Dm}{k}\right\rfloor\frac{d(\delta,1-\delta)}{d(1-p,p)}\geq\left\lfloor\frac{E}{x}\right\rfloor\left\lfloor\frac{D}{k}\right\rfloor\cdot mn\cdot\frac{d(\delta,1-\delta)}{d(1-p,p)}.$ (31)
@@
$\sum_{C_{1,k}}\sum_{V_{4,x}}\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\leq\sum_{i\in V}\sum_{j\in C}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]=\operatorname{\mathbb{E}}^{\nu}\left[\sum_{i\in V}\sum_{j\in C}T_{(i,j)}\right]=\operatorname{\mathbb{E}}^{\nu}[\tau],$ (32)

### v2

@@ Appendix C C   Lower bound for non-adaptive algorithms in the incomplete information setting
We denote a given non-adaptive committee selection algorithm by $\operatorname{\textsc{alg}}$. We define the class of instances that will serve as an example for the lower bound: let $k=2p+r$, where $p\in\operatorname{\mathbb{N}}$ and $r\in\{0,1\}$, which means we can form any $k\in\operatorname{\mathbb{N}}_{\geq 2}$ for the committee size. Partition the candidates into $C_{1}\cup C_{2}\cup\ldots\cup C_{p}\cup D$, with $|C_{i}|=\lfloor(m-r)/p\rfloor$ for all $i\in\{1,\ldots,p\}$, $D$ containing the remaining candidates. Because $\sum_{i=1}^{p}|C_{i}|=p\cdot\lfloor(m-r)/p\rfloor\leq m-r$, we know $|D|\geq r$. For each $i$, we form $S_{i}=\{c_{i,1},\ldots,c_{i,\ell}\}\subseteq C_{i}$ by picking $\ell$ distinct candidates from $C_{i}$. The candidates are all chosen to join $S_{i}$ with the same probability, independently of one another, and the assignment of these candidates to the numbers $1,\ldots,\ell$ is also random. Moreover, this procedure happens independently for all $C_{i}$. We partition the voters into $p$ parties $P_{1},\ldots,P_{p}$ and, if $r=1$, one party $R$ (if $r=0$ there is no set $R$). Each party $P_{i}$ contains $2/k$ of the voters and all voters in $P_{i}$ approve only of candidates within $S_{i}$ (so no one approves any candidate in $C_{i}\setminus S_{i}$). More specifically, for each $P_{i}$, we can again partition the candidates of $C_{i}$ into two groups with non-overlapping approval sets: candidates $c_{i,1},\ldots,c_{i,\ell-1}$, are collectively approved by $1/2$ of the voters in $P_{i}$ (so $1/(2k)$ voters in total) and $c_{i,\ell}$, that we call the designated candidate, is approved by the other half of $P_{i}$. The party $R$, if it exists, contains $1/k$ of the voters, all of whom approve precisely one candidate $d\in D$.
We say that $\operatorname{\textsc{alg}}$ covers a set of candidates $S$ if $\operatorname{\textsc{alg}}$ ever submits a query set $Q$ for which $|Q\cap S|\geq 2$. This means that, if for any of the parties $P_{i}$, $\operatorname{\textsc{alg}}$ fails to cover $S_{i}$, all $\ell$ of these candidates are indistinguishable to $\operatorname{\textsc{alg}}$, given the definition of our query model. Indeed, all candidates within $S_{i}$ are approved by the same number of voters, and we can only start to distinguish the structure of $S_{i}$ if we combine $c_{i,\ell}$ and $c_{i,j}$ for some $j\in\{1,\ldots,\ell-1\}$ in one query. Then, if a set $S_{i}$ is not covered, the best $\operatorname{\textsc{alg}}$ can do is pick candidates from $S_{i}$ randomly, since without covering $S_{i}$ it is still possible to distinguish which elements form the set $S_{i}$.
For each $i$, writing $k_{i}$ for the number of candidates $\operatorname{\textsc{alg}}$ selects from $C_{i}$, $\operatorname{\textsc{alg}}$ picks the designated candidate with probability at most $k_{i}/\ell$. Suppose that for all $i$, $\operatorname{\textsc{alg}}$ picks some candidate $c_{i,j}$ for $j\in\{1,\ldots,\ell-1\}$ and also picks the single approved candidate in $R$. We are allowed to assume this since this assumption facilitates $\operatorname{\textsc{alg}}$. This already gives a Max Cover score of $\frac{p+1}{2p+1}$. Then, in order to attain a $\gamma(1-1/\operatorname{\text{e}})$-max cover score, $\operatorname{\textsc{alg}}$ must still pick $\gamma(1-1/\operatorname{\text{e}})(2p+1)-(p+1)$ designated candidates. Then if $t=1$, it is impossible for $\operatorname{\textsc{alg}}$ to cover any of the $S_{i}$ and we attain a $\gamma(1-1/\operatorname{\text{e}})$-max cover ratio only if $\operatorname{\textsc{alg}}$ is lucky enough to select at least $\left\lceil\left(1-\frac{1}{\operatorname{\text{e}}}\right)\gamma+\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)2\gamma-1\right)p-1\right\rceil=:q$ of the designated candidates without having any information about which ones to pick. Clearly, for each $C_{i}$, the probability of picking the correct candidate is at most $p/\ell$, as $p$ is the total number of candidates we can still pick, and $\ell$ is the size of each set. The probability to pick at least $q\geq 1$ designated candidates is at most the probability to pick a single one: $p/\ell$.
Now suppose that $t>1$ and so we might be able to cover some $S_{i}$. Moreover, suppose that $\operatorname{\textsc{alg}}$ knows the partition of candidates into $C_{1}\cup\ldots\cup C_{p}\cup D$ (so it will query within these sets) and knows the distribution of approvals within the sets $C_{i}$ (but it doesn’t know how this corresponds to the particular candidates). Moreover, we assume we can make at most $U\cdot m^{2}$ queries within each set $C_{i}$, for some constant $U$.
@@
sets of size $\ell$ in $C_{i}$ in total, and since all sets were chosen to join $S_{i}$ with the same probability, $\operatorname{\textsc{alg}}$ covered $S_{i}$ with probability at most
Suppose that $\operatorname{\textsc{alg}}$ picked at least one of the candidates $c_{i,j}$ for $j\in\{1,\ldots,\ell-1\}$ for all $i$, then, like before, we need to pick at least $q$ designated candidates in order to attain the desired ratio. $\operatorname{\textsc{alg}}$ can do this by covering no $S_{i}$ and picking all $q$ designated candidates by chance, which happens, like proven above, with probability at most $p/l$. $\operatorname{\textsc{alg}}$ can also cover one of the sets $S_{i}$ and pick q1q-1 designated candidates by chance, or any other combination is possible up to covering all sets $S_{i}$. Then, for $t>1$,
@@
where the inequality is due to the fact that covering a set $S_{i}$ is necessary to distinguish its designated candidate, but might not be sufficient. For $h=0$, we have bounded the probability above as $p/\ell$. Each of the other at most (pq)p(p-q)p probabilities is bounded from above by the probability of covering one set $S_{i}$, which is at most $\ell!2^{\ell+t}Up^{2}$. If we fix $U=\frac{\delta}{\ell!2^{t+\ell}p^{3}(p-q)}$ this becomes at most $\delta$. The total probability is then at most $\frac{p}{\ell}+\delta$. Recall that we want this to be at most $\epsilon$, which we can achieve, for instance, by fixing both terms to $\epsilon/2$. Since $\delta$ determines the query complexity, we can choose $\delta$, so it suffices to show that $\frac{p}{\ell}\leq\frac{\epsilon}{2}$. Since p=2k1p=2k-1, this condition is met when $l\geq\frac{k-1}{\epsilon}$.
As mentioned above, we have at most $Um^{2}$ queries and $U=\frac{\epsilon/2}{\ell!2^{t+\ell}p^{3}(p-q)}$, where p=(k1)/2p=(k-1)/2. We want to denote the scaling of the complexity in terms of $\gamma$ and $\epsilon$. Since p=(k1)/2p=(k-1)/2 and $q=\left\lceil\left(1-\frac{1}{\operatorname{\text{e}}}\right)\gamma+\left(\left(1-\frac{1}{\operatorname{\text{e}}}\right)2\gamma-1\right)p-1\right\rceil$, we obtain
@@
where the last step required $2k\gamma/2\geq 1/2$, which is true in case $k\geq 2,\gamma>1/2$. It is not hard to see that $\frac{1}{\ell!2^{t+\ell}p^{3}}\leq 1$, so that the final query complexity becomes $\frac{m^{2}\epsilon}{1-\gamma}$. ∎
Theorem 14 motivates the study of adaptive algorithms. It might be possible to extend this result to a larger exponent of $m$, as done in [14]), by adapting the instance. We leave this for future work.
Appendix D D   Inaccurate information: proofs of Theorem 7 and Theorem 8
We prove the upper and lower bound on achieving the optimal approximation ratio for Max Cover in the $p$-inaccurate query model. We stated these results as Theorem 7 and Theorem 8, respectively.
Recall that for accurate information, we write $A(i,j)$ to query the opinion of voter $v_{i}\in V$ on candidate $c_{j}\in C$, where the outcome is 1 in case of approval and 0 in case of disapproval. In the $p$-inaccurate query model, we instead query $A_{p}(i,j)=A(i,j)\oplus X$ where $X\sim\text{Bernoulli}(p)$, so that the outcome is flipped with probability $p$. This error is added independently for each pair $(i,j)$ and each repeated sampling. For the upper bound, we give an algorithm which recovers the optimal approximation ratio with high probability. Our goal is to minimize the number of queries to $A_{p}(i,j)$.
@@ Appendix D D   Inaccurate information: proofs of Theorem 7 and Theorem 8
For all $v_{i}\in V,\,c_{j}\in C$, perform query $A_{p}(i,j)$ $U=\left\lceil 2\frac{\log((nm)/\delta)}{1/\log(4p(1-p))}\right\rceil$ times and take the majority winner as the ‘true’ query response;
Let $W=\{\}$ be an empty set;
@@
    For all $c\notin W$, determine $\Delta(W,c)$.
@@
    Add to $W$ any $c^{\prime}\in\operatorname*{argmax}_{c\notin W}\,\Delta(W,c)$.
@@
Theorem 7.
Let $p\in(0,\frac{1}{2})$, $\delta>0$, $n,m\in\mathbb{N}$. Then there exists an algorithm that is (11/e)(1-1/e)-approximate for Max Cover in the $p$-inaccurate query model w.p. at least $1-\delta$ and with query complexity $O(nm\log(nm/\delta))$.
@@
We adapt Algorithm 1, obtaining Algorithm 6. While Algorithm 1 simply queries $A(i,j)$ once for each of the $nm$ voter-candidate pairs, Algorithm 6 instead performs each query $A_{p}(i,j)$ some $U>1$ number of times and takes the majority winner to be the ‘true’ value, breaking ties arbitrarily. We now show that, taking $U$ as in Algorithm 6, each of the $nm$ majority votes are correct with probability at least $1-\frac{\delta}{nm}$. Then, by independence, all together are correct with probability $1-\delta$, in which case we obtain a $(1-1/\operatorname{\text{e}})$-approximate solution. Each majority vote fails if a strict majority of the $U$ samples is incorrect, and we can bound this failure probability using a Chernoff bound: Take $\Omega=\{0,1\}^{U}$, $\mathcal{F}=\mathcal{P}(\Omega)$ and $\mathbb{P}:\mathcal{F}\to[0,1]$ that satisfies $\mathbb{P}(\{\mathbf{v}\})=(1-p)^{y}p^{U-y}$ when $\mathbf{v}$ contains $y$ 1’s and UyU-y 0’s. Accordingly, write $\mathbf{v}_{i,j}$ for a vector that is the outcome of $U$ throws of $(i,j)$, where entry 0 means the outcome is corrupted and 1 means the outcome is correct. Then take $X:\Omega\to\operatorname{\mathbb{N}}:\mathbf{v}_{i,j}\mapsto\sum_{l=1}^{U}v_{i,j,l}$ for the random variable that maps a vector to the number of correct outcomes and write $\bar{X}:=U-X$ for the number of failures. A Chernoff bound then tells us that $\mathbb{P}(\bar{X}\geq\frac{1}{2}U)\leq\exp(-UD(\frac{1}{2}||p))$ [13] and we can reformulate to
$\begin{split}\exp\left\{-UD\left(\frac{1}{2}||p\right)\right\}&:=\exp\left\{-U\left(\frac{1}{2}\log\left(\frac{1/2}{1-p}\right)+\frac{1}{2}\log\left(\frac{1/2}{p}\right)\right)\right\}\\ &=\exp\left\{-U\frac{1}{2}\log\left(\frac{1/2}{1-p}\right)\right\}\exp\left\{-U\frac{1}{2}\log\left(\frac{1/2}{p}\right)\right\}\\ &=\left(\frac{1}{4(1-p)p}\right)^{-\frac{1}{2}U}.\end{split}$ (23)
We determine the value of $U$ so that the above is at most $\frac{\delta}{nm}$:
$\left(\frac{1}{4(1-p)p}\right)^{-\frac{1}{2}U}\leq\frac{\delta}{nm}\iff-\frac{1}{2}U\leq\log_{\frac{1}{4p(1-p)}}\left(\frac{\delta}{nm}\right)$
$\iff U\geq-2\log_{\frac{1}{4p(1-p)}}\left(\frac{\delta}{nm}\right)=-2\frac{\log\frac{\delta}{nm}}{\log\frac{1}{4p(1-p)}}=2\frac{\log\frac{nm}{\delta}}{\log\frac{1}{4p(1-p)}}.$
Fixing $\left\lceil 2\frac{\log((nm)/\delta)}{1/\log(4p(1-p))}\right\rceil$ thus guarantees that all estimates are correct, and thus the approximation ratio is achieved with probability at least $1-\delta$. ∎
Although the above proof involved an adaptation of the Greedy algorithm, we note that we could have just as easily adapted the local search algorithm. Both algorithms will query all $n$ voters on all $m$ candidates. To deal with the $p$-inaccurate information setting, the above proof simply shows that repeating each of the $nm$ queries $U=\log(nm/\delta)$ times guarantees that the $nm$ queries are all correct with probability $1-\delta$. This guarantee thus holds independently of what the algorithm does with the query outcomes, and also holds for the local search algorithm. This also implies that the above upper bound holds for the generalized problem of optimizing over a matroid contraint.
Theorem 8.
@@
Let $p\in(0,\frac{1}{2})$, $\delta>0$, $n,m\in\mathbb{N}$. Then any algorithm that is (11/e)(1-1/e)-approximate for Max Cover in the $p$-inaccurate query model w.p. at least $1-\delta$ has expected query complexity $\Omega(nm\log(1/\delta))$.
@@
To establish the theorem, we adapt the proof of Theorem 9 in [19], using a result from multi-armed bandit theory [20] to lower bound the query complexity. The Kullback-Leibler divergence between two probability distributions is defined as $D(P||Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}$ for discrete probability functions $P$ and $Q$, as long as, for all $x,\,Q(x)=0\implies P(x)=0$ (otherwise we define it as $\infty$). This reduces to $d(a,b)=a\log\frac{a}{b}+(1-a)\log\frac{1-a}{1-b}$ for two Bernoulli distributions with parameters $a,b\in(0,1)$.
@@
Recall that we query voter $v_{i}\in V$ on candidate $c_{j}\in C$ using $A_{p}(i,j)=A(i,j)\oplus X$ where $X\sim Bernoulli(p)$. We will view each query $A_{p}(i,j)$ as an arm $a\in\mathcal{A}$ in a multi-armed bandit problem, and write $y_{t}:=A_{p}(i,j)_{t}$ for the outcome of the $t$-th query. When $A(i,j)=0$. Then $A_{p}(i,j)\sim\text{Bernoulli}(p)$. Conversely, when $A(i,j)=1$ then $A_{p}(i,j)=\text{Bernoulli}(1-p)$. We let $P_{a}$ denote the Bernoulli distribution associated with action $a$: for any query index $t$ and any $y\in\{0,1\}$ we have $P_{a}(y)=\mathbb{P}(y_{t}=y|a_{t}=a)$. Let us refer to an arbitrary (possibly randomized) algorithm for the problem by $\operatorname{\textsc{alg}}$. We write $\operatorname{\mathbb{E}}^{\nu}$ and $\mathbb{P}^{\nu}$ for the expectation and probability when the underlying instance is $\nu$. Write $L_{\nu}$ for the set of optimal committees for instance $\nu$. By assumption we have $\mathbb{P}^{\nu}(\operatorname{\textsc{alg}}(\nu)\in L_{\nu})\geq 1-\delta$ for any instance $\nu$. We denote the random variable that portrays the number of queries taken when $\operatorname{\textsc{alg}}$ terminates, by $\tau$. We use the following lemma from multi-armed bandit theory, that is mentioned e.g. in [20], Chapter 15 (Lemma 15.1 and more specifically Exercise 15.7).
Lemma 15.
@@
For any $k\geq 2$, consider an instance $\nu$ with $m$ candidates and $n$ voters. We partition the voters into $m$ disjoint “parties” $P_{1},\ldots,P_{m}$. Voters in party $P_{i}$ approve precisely $c_{i}$. Candidates $c_{1},\ldots,c_{m-k}$ (together $C_{1}$) all have one approving voter. These mkm-k voters (together $V_{1}$) are alone in their party. Candidates $c_{m-k+1},\ldots,c_{m}$ (together $C_{2}$) all have $u>\frac{1}{1-\frac{1}{\operatorname{\text{e}}}\cdot k}$ approving voters (together $V_{2}$), so that the only $(1-\frac{1}{\operatorname{\text{e}}})$-approximate solution consists of the $k$ candidates in $C_{2}$:
@@
This shows even holding in the committee k1k-1 candidates from $C_{2}$ and one candidate from $C_{1}$, instead of $k$ candidates from $C_{2}$, yields an unforgivable loss in score. Finally we have $V_{3}$, a set of voters that approve of no candidate (homeless voters). For any $m,n$, we can find constants $D,E\in(0,1)$ such that $|C_{1}|\geq D\cdot m$ (which means we need $0\leq m-Dm-k$) and $(1-E)n-ku\geq|C_{1}|$ so that $|V_{3}|\geq E\cdot n$, so let’s pick any such constants $D$ and $E$. Now consider the instance $\nu^{\prime}$ that is obtained by taking, $x:=\lceil k(\frac{u}{1-1/\operatorname{\text{e}}}-1)\rceil$ homeless voters and distributing them (as) evenly (as possible) over $k$ candidates $c_{1},\ldots,c_{k}\in C_{1}$. Other than that, all preferences remain unchanged. With this value of $x$, we have $x+k>\frac{ku}{1-1/\operatorname{\text{e}}}$, which means that $\nu^{\prime}$ has a unique $(1-1/\operatorname{\text{e}})$-approximate solution in $c_{1},\ldots,c_{k}\subseteq C_{1}$, whereas the unique $(1-1/\operatorname{\text{e}})$-approximate solution of $\nu$ was the set $C_{2}$. Since both instances have only one optimal solution, and these differ, we have $L_{\nu}\cap L_{\nu^{\prime}}=\emptyset$.
@@
Now, regarding the inequality in Lemma 15, note the following. Firstly, $D(P_{a}||P^{\prime}_{a})=0$ for all $a\notin\{(i,j):i\in\{1,\ldots,x\},j\in\{1,\ldots,k\}\}$. When $D(P_{a},P^{\prime}_{a})\neq 0$, it equals d(1p,p)d(1-p,p). Let $\mathcal{E}$ be the event that $\operatorname{\textsc{alg}}$ outputs (the only element in) $L_{\nu^{\prime}}$. Then, by assumption, $\mathbb{P}^{\nu}(\mathcal{E})\leq\delta$ and $\mathbb{P}^{\nu^{\prime}}(\mathcal{E})\geq 1-\delta$. Monotonicity of the discrete KL function $d$ implies that $d(\mathbb{P}^{\nu}(\mathcal{E}),\mathbb{P}^{\nu^{\prime}}(\mathcal{E}))\geq d(\delta,1-\delta)$.
So now we can rewrite the inequality in Lemma 15 as
@@
$\sum_{C_{1,k}}\sum_{V_{3,x}}\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\geq\left\lfloor\frac{En}{x}\right\rfloor\left\lfloor\frac{Dm}{k}\right\rfloor\frac{d(\delta,1-\delta)}{d(1-p,p)}\geq\left\lfloor\frac{E}{x}\right\rfloor\left\lfloor\frac{D}{k}\right\rfloor\cdot mn\cdot\frac{d(\delta,1-\delta)}{d(1-p,p)}.$ (27)
@@
$\sum_{C_{1,k}}\sum_{V_{3,x}}\left(\sum_{i=1}^{x}\sum_{j=1}^{k}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]\right)\leq\sum_{v_{i}\in V}\sum_{c_{j}\in C}\operatorname{\mathbb{E}}^{\nu}[T_{(i,j)}]=\operatorname{\mathbb{E}}^{\nu}\left[\sum_{v_{i}\in V}\sum_{c_{j}\in C}T_{(i,j)}\right]=\operatorname{\mathbb{E}}^{\nu}[\tau],$ (28)